OpenStudy (anonymous):

A triangular playground has angles whose measures are in the ratio 8:3:9. What is the measure of the smallest angle?

OpenStudy (anonymous):

the sum of the angles of a trianlge is 180 given the ratio, we can formulate the formula 8x+3x+9x=180 ; where x is the constant of proportionality of the angles. 8x+3x+9x = 180 20x = 180 x = 9 therefor, the smallest angle 3x = 3(9) = 27 degrees